im looking for the answer to f(g(x)) and g(f(x)), when f(x)=10x+1 and g(x)=sqrt 2x+1
Answer by stanbon(48535)

g(f(x))= g(10x+1)= sqrt[2(10x+1)+1]= sqrt[20x+3]
f(g(x)) = f(sqrt(2x+1))= 10[sqrt(2x+1)]+1
Cheers,
Stan H.